XY2 dissociates as, XY2(g)= XY(g) + Y(g) (reversible reacyion)..Initial pressure of XY2 is 600 mm Hg.THe total pressure at equlibrium is 800 mm hg.Assuming volume of system  to reamin constant ,the value of Kp is.?

Grade:12

4 Answers

Naveen Kumar
9 years ago
XY2(g).............= …...........XY(g) + Y(g)
600mm.............................0...........0......(initially)
600-p..................................p..........p....(at equilibrium)
[Here I have used directely the partial pressure because as the volume is constant, so number of moles of reqactant or product∝ partial pressure]
Now applying Dalton’s law of partial ppressure at equilibrium,( total pressure=sum of partial pressure of all the gas components in the mixture)
800=(600-p) + p+ p=600+p
p=200mm of Hg
writing Kp=p*p/(600-p);
put value of p and change the unit according to :105 Pa=760mm of Hg
Ramreddy IIT Patna
9 years ago
Kp= (pXY) x (pY)/ (pXY2)
XY2 ---------> XY + Y
600 mm 0 0
600 -x x x
Total pressure at equillibrium = 600+x = 800
x= 200
Kp= (200 x 200) / 600
..............................................................................................................
Sahil Saha
6 years ago
XY2=XY+Y
[XY2] at equilibrium=600 atm
Total Pressure= 800 atm
Kp=[XY][Y]/[XY2]
[XY]=[Y]=x (let)
Total pressure at equilibrium = 600+x=800
x=200
Kp=200x200/600
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your question.
 
Kp= (pXY) x (pY)/ (pXY2)
XY2 ---------> XY + Y
600 mm       0      0
600 – x        x      x
Total pressure at equillibrium = 600 + x = 800
x = 200
Kp = (200 x 200) / 600 = 66.67 mm = 8772 Pa
 
Thanks and regards,
Kushagra